OpenStudy (anonymous):

Check my answers for these word problems? 1. A car rents for $18.75 per day, plus $.25 per mile. If you rent the car for one week, what is the greatest number of miles you can drive and still keep the cost of renting the car under $200 for the week? my answer: 274 miles 2. The sum of twice a number and –6 is nine more that the opposite of the number. Find the number. my answer: 15 Thanks for checking! :)

OpenStudy (anonymous):

for 2nd question i think it shoulbe 2x-6=9-2x

OpenStudy (anonymous):

okay thanks! I knew the second one was wrong, but was the first one right?

OpenStudy (anonymous):

for the second one is the number 3.75?

OpenStudy (anonymous):

sorry 137.5 for 1st one

OpenStudy (anonymous):

for the first one i multiplied 18.75 times 7 b/c they rent for one week= 7 days equals $131.25 then if it is .25 for each mile, times 274 = $68.5 add 68.5 + 131.25 = 199.75 under $200 .....so how is my first answer wrong?

OpenStudy (anonymous):

oh sorry i took 0.5 instead of 0.25 yes u are correct :)

OpenStudy (anonymous):

thank you for all your help! (:

OpenStudy (anonymous):

pleasure!